OpenStudy (anonymous):

Circle O has a diameter . If A is at (-6, -2) and B is at (-3, 4), find the center of the circle.

OpenStudy (anonymous):

a.)3/5,-3 b.)-9/2,1 c.)5,-6 d.)3,6

OpenStudy (anonymous):

\[(\frac{ x1+x2}{ 2 },\frac{ y1+y2 }{ 2 })=(c1,c2)\] formula substitute

OpenStudy (anonymous):

(x1,y1)=(-6,-2),(y1,y2)=(-3,4) center =(c1,c2 )

OpenStudy (anonymous):

option B hope u got it .....

OpenStudy (anonymous):

i did thanks

OpenStudy (anonymous):

u welcome:-)))

OpenStudy (anonymous):

wanna help me somemore??